OpenStudy (anonymous):

Which of the following is a factor for xy^2n^2 + x^2Y^2n: I know the answer is xy^2n but i do not know how to get there

OpenStudy (anonymous):

because \(xy^2n\times n=xy^2n^2\) and \(xy^2n\times x=x^2y^2n\)